Calculate the uncertainty of the velocity of an electron, if the position uncertainty is 2.52×10^-11 m.
1
Expert's answer
2019-09-26T05:53:41-0400
"\Delta\chi\Delta p_x\geqslant h"

"\Delta\chi m_{e^-}\Delta v = h"

Where "m_{e^-} = 9.109*10^{-31}kg, h = 6.626*10^{-34}J*sec"

Then

"\Delta v = {h \over \Delta\chi m_{e^-}} = {6.626*10^{-34}J*sec \over 2.52*10^{-11}m*9.109*10^{-31}kg} = 0.289*10^8 m/sec = 2.89*10^7m/sec"

Answer: 2.89*107 m/sec.